OpenStudy (anonymous):

(m - 18 / 5) - (4m - 14 / 7) = 1 I need help solving this

OpenStudy (anonymous):

Solving it for M?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

m - 18/5 - 4m + 2 = 1.

OpenStudy (anonymous):

-3m = 1 - 2 + 18/5

OpenStudy (anonymous):

How did you get + 2

OpenStudy (anonymous):

Subtracting a negative number gives you a positive. Just like multiplying two negatives gives you a positive value.

OpenStudy (anonymous):

The way I solved it: With the LCD 35 35(m - 18 / 5) - 35(4m - 14 / 7) = 35(1) (35m - 18 / 5) - 140m - 14 / 7) = 35 (35m - 18 / 5) - (140m - 14 / 7) = 35 ..

OpenStudy (anonymous):

@wouterdobber

OpenStudy (anonymous):

You are not distributing into both values inside the parenthesis with the lcd

OpenStudy (anonymous):

Exactly what rachelfreak said.

OpenStudy (anonymous):

But I do not think it's necessary to multiply by 35 though.

OpenStudy (anonymous):

i agree with wouterdobber

OpenStudy (anonymous):

Where we got to before: -3m = 1 - 2 + 18/5. From here you could just make it, -5/5 +18/5 = -3M. So: 13/5 = -3M So: 65/5 = -M So 15 = -M So M = -15

OpenStudy (anonymous):

Wait I messed up.

OpenStudy (anonymous):

39/5 = -M So: M = - 39/5 So M = - 7.8 I think.